Determine the surface area of the portion of 2x + 3y + 6z = 9 that is in the 1st octant.


1
Expert's answer
2021-05-07T12:42:12-0400

We have given the portion,


3x + 2y + z = 9


The plane intersects the first octant in a triangle with the vertices "({2},0,0),(0,3,0),(0,0,6)"


since these are the intercepts with the positive x, y and z axes respectively. Thus this is the


surface area of the part of the surface "z= 6-3x-2y" over the region "0 \le x \le {2}" , "0 \le y\le 3 - \dfrac{3x}{2}"


The derivatives of the function "6-3x-2y" in the x and y directions are -3 and 2, and so the surface area is


"\int_{0}^{2} \int_{0}^{3-\dfrac{3x}{2}} \sqrt{1+(-3)^2+(-2)^2}dydx = 3 \sqrt{14}"